OpenStudy (anonymous):

A school has five different after-school activities planned in the fall. Kim has time to participate in two of these activities. How many different pairs of after-school activities can Kim chose from the available activities?

ganeshie8 (ganeshie8):

hint : number of ways of choosing `k` objects from a collection of `n` objects is \[\large \binom{n}{k} \]

ganeshie8 (ganeshie8):

hint2 : \[\large \binom{n}{k} = \dfrac{n!}{k!(n-k)!} \]

OpenStudy (anonymous):

sorry i am going to try to figure this out really quick haha! i will tag you with the answer i get if that is ok? @ganeshie8

ganeshie8 (ganeshie8):

sure :) you're familiar with factorials right ?

OpenStudy (anonymous):

yes i am :)

ganeshie8 (ganeshie8):

great ! then just plug the numbers and simplify : n = 5 k = 2

ganeshie8 (ganeshie8):

number of ways of choosing `2` objects from a collection of `5` objects is \[\large \binom{5}{2} = \dfrac{5!}{2!(5-2)!} \] \[\large = ?\]

OpenStudy (anonymous):

ok i got 20??

OpenStudy (anonymous):

^^ @ganeshie8

ganeshie8 (ganeshie8):

very close, but no - looks you forgot to divide by 2

OpenStudy (anonymous):

oh sorry forgot one of the factorials hold on let me work this out

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

ok i got 10 this time sorry about that :)

ganeshie8 (ganeshie8):

10 is \(\large \color{red}{\checkmark }\)

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

Thank you so much

ganeshie8 (ganeshie8):

you could have tried this on your own without using the fancy formula as well

ganeshie8 (ganeshie8):

wana see how to work this out with out using the formula ?

OpenStudy (anonymous):

yeah sure!

OpenStudy (anonymous):

also is the formula for a permutation the same as a combination

ganeshie8 (ganeshie8):

they look similar, but permutation formula will not have one factorial in the denominator : \[\large ^nP_r = \dfrac{n!}{(n-r)!}\]

ganeshie8 (ganeshie8):

notice that there is not \(\large r!\) in the denominator

OpenStudy (anonymous):

oh thank you

ganeshie8 (ganeshie8):

check this link when you're free http://www.mathsisfun.com/combinatorics/combinations-permutations.html it has excellent explanation for both permutations and combinations

OpenStudy (anonymous):

thatk you again this was so helpful

ganeshie8 (ganeshie8):

np :)